OpenStudy (anonymous):

Omar is 3 times as old as William and is also 18 years older than William. How old is William?

OpenStudy (anonymous):

How can William be 18 years older than William?

OpenStudy (anonymous):

^^^

OpenStudy (anonymous):

Lol these questions are getting to me :p I don't know why but I have a soft spot for this kid!

OpenStudy (anonymous):

9?

OpenStudy (anonymous):

Am I right, can other people check it.

OpenStudy (anonymous):

9*3=27 27-18=9

OpenStudy (raffle_snaffle):

^did you set them equal

OpenStudy (anonymous):

Yeah he's 9

OpenStudy (anonymous):

oh I thought it was 6 bc 18/3=6

OpenStudy (raffle_snaffle):